# Compatible Poisson Structures On Multiplicative Quiver Varieties

### v2

@@ 3.3.1. Deformation of the moment map
$P^{\gamma}:=$ $\,\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\gamma_{a}\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})\,\partial_{a}\wedge\partial_{a^{\ast}}\right]$ (3.16)
@@
$\omega^{\gamma}:=$ $\,-\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\gamma_{a}\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}\,\mathrm{d}\mathtt{X}_{a}\wedge\mathrm{d}\mathtt{X}_{a^{\ast}}\right]$ (3.17)
@@ 3.3.2. Character varieties
The smooth complex variety $M_{g,r,n}$ is endowed with a Hamiltonian quasi-Poisson pencil of order $g(g-1)/2$ centered at the quasi-Poisson bivector
$P_{\mathrm{char}}:=$ $\,\frac{1}{2}\sum_{1\leq i\leq g}\,\operatorname{tr}\left[A_{i}^{\ast}A_{i}\,\partial_{A_{i}}\wedge\partial_{A_{i}^{\ast}}-A_{i}A_{i}^{\ast}\partial_{A_{i}^{\ast}}\wedge\partial_{A_{i}}\right]+\frac{1}{2}\sum_{1\leq j\leq r}\operatorname{tr}\left[Z_{j}^{2}\,\partial_{Z_{j}}\wedge\partial_{Z_{j}}\right]$ (3.19)
@@ 3.3.3. Generalized MQV
[…]
@@
[…]
@@ 3.3.4. Quiver varieties
It was suggested by A. Alekseev that the Poisson structures defined in the previous two results can be obtained as special instances of a construction due to Lu and Mouquin [30]. Namely, for Proposition 3.19, see $G$ as a Poisson-Lie group for the zero Poisson structure, so that $\mathfrak{g}=\mathfrak{g}^{\prime}\oplus\bigoplus_{j=1}^{\ell}\mathfrak{g}_{j}$ is a Lie bialgebra with cobracket $\delta_{\mathfrak{g}}$ being trivially the sum of the zero cobrackets on $\mathfrak{g}^{\prime},\mathfrak{g}_{1},\ldots,\mathfrak{g}_{\ell}$. If $1^{(j)}\in\mathfrak{g}$ denotes the image of $1\in\mathbb{C}$ induced at the level of Lie algebras by the inclusions ×$\mathbb{C}^{\times}\hookrightarrow G_{j}\hookrightarrow G$, then $\mathtt{t}=\sum_{i<j}z_{ij}1^{(i)}\wedge 1^{(j)}\in\wedge^{2}\mathfrak{g}$ is a mixed twisting element [30, Def. 3.1]. It follows from [30, Lem. 2.6] that $(M,P+\mathtt{t}_{M})$ is a Poisson $(\mathfrak{g},\delta_{\mathfrak{g},\mathtt{t}})$-variety for the twisting cobracket $\delta_{\mathfrak{g},\mathtt{t}}=\delta_{\mathfrak{g}}+[\mathtt{t},-]$; but the latter vanishes while the action of $\mathfrak{g}$ on $M$ sends $\mathtt{t}$ to $\mathtt{t}_{M}=\psi_{\underline{z}}$ (3.4), hence $(M,P+\psi_{\underline{z}})$ is a Poisson $G$-variety.
As for Theorem 3.20, consider for each $a\in Q$ the subquiver $\Gamma_{a}=t(a)\underset{a^{\ast}}{\stackrel{{\scriptstyle a}}{{\rightleftarrows}}}h(a)$ of $\overline{Q}$. We have an action of $G_{a}:=\operatorname{GL}_{n_{t(a)}}\times\operatorname{GL}_{n_{h(a)}}$ on $Y_{a}:=M_{\Gamma_{a}}(n_{t(a)},n_{h(a)})$, cf. (2.10), thus $M_{\overline{Q}}=\times_{a\in Q}Y_{a}$ inherits an action of ×$\times_{a\in Q}G_{a}$ (it is “before fusion”, see Remark 2.11). Denote the Lie algebra of $G_{a}$ by $\mathfrak{g}_{a}$, and see $\mathfrak{g}=\oplus_{a\in Q}\mathfrak{g}_{a}$ as a Lie bialgebra for the zero cobracket as above. Let $1^{(a)}$ be the image of $(\operatorname{Id}_{n_{t(a)}},0_{n_{h(a)}})$ under the inclusion $\mathfrak{g}_{a}\hookrightarrow\mathfrak{g}$. Again, $\mathtt{t}=\sum_{a<b}z_{ab}1^{(a)}\wedge 1^{(b)}$ is mixed twisting and induces $\psi_{\underline{z}}$ (3.13). Hence $P_{\mathtt{qv}}+\psi_{\underline{z}}$ is a mixed product Poisson structure on $M_{\overline{Q}}$ for the action of ×$\times_{a\in Q}G_{a}$, cf. [30, Prop. 3.4]. Factoring the action of $\operatorname{GL}_{\mathbf{n}}$ through ×$\times_{a\in Q}G_{a}$ by taking diagonal embeddings yields the Poisson $\operatorname{GL}_{\mathbf{n}}$-variety structure considered in Theorem 3.20.
@@ 4. Noncommutative origin of the pencil
Fix an associative unital $\mathbb{C}$-algebra $A$, and consider $A\otimes A$ as a $\mathbb{C}$-algebra for the product satisfying $(f_{1}\otimes f_{2})(f_{3}\otimes f_{4})=f_{1}f_{3}\otimes f_{2}f_{4}$, $f_{1},\ldots,f_{4}\in A$. Denote by $\tau_{(12)}\in\operatorname{End}(A\otimes A)$ the operation that exchanges the two tensor factors. An outer double bracket [^9] This is simply called a double bracket by Van den Bergh [39]. We dub this operation as outer because the derivation rule of the second argument is defined thanks to the outer $A$-bimodule structure on $A\otimes A$ which is determined by $f_{1}\ast(g\otimes h)\ast f_{2}=f_{1}g\otimes hf_{2}$ for any $f_{1},f_{2},g,h\in A$. Below, we shall use right double brackets [20] based on the right $A$-bimodule structure $f_{1}\cdot_{r}(g\otimes h)\cdot_{r}f_{2}=g\otimes f_{1}hf_{2}$; hence qualifying the type of double bracket removes any possible confusion. is a bilinear map {{,}}$\left\{\!\!\left\{-,-\right\}\!\!\right\}_{o}:A\times A\to A\otimes A$ such that for $f,g,h\in A$,
@@
An important result is that the pair $(A_{Q},\left\{\!\!\left\{-,-\right\}\!\!\right\}_{o})$ is a quasi-Poisson algebra [15, Theorem 3.3]. Without unravelling the definition (cf. [39, §5.1]) let us only give consequences of this fact. Put $H_{0}(A_{Q})=A_{Q}/[A_{Q},A_{Q}]$, which is a vector space coming with a linear map $\operatorname{tr}:A_{Q}\to H_{0}(A_{Q})$ that sends commutators to zero. We get from [39, Lem. 5.1.3] that $H_{0}(A_{Q})$ is equipped with a Lie bracket, denoted {,}$\{-,-\}_{\mathtt{VdB}}$, uniquely determined by
@@
where $\mathrm{m}:A_{Q}\otimes A_{Q}\to A_{Q}$ is the multiplication. In particular, (4.3) extends to a Poisson bracket on $\operatorname{Sym}(H_{0}(A_{Q}))$, which we also denote by {,}$\{-,-\}_{\mathtt{VdB}}$. The following result follows from [39, §7.7,7.13] and the fact that the representation space $\operatorname{Rep}(A_{Q},\mathbf{n})$ coincides with $M_{\overline{Q}}^{\gamma}(\mathbf{n})$ for any $\mathbf{n}\in\mathbb{Z}^{I}_{\geq 0}$.
@@
is a morphism of Poisson algebras if $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with {,}$\{-,-\}_{\mathtt{VdB}}$ (4.3) and $\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ is equipped with $P^{\gamma}$ (3.16).
@@
To understand the bivector $\psi_{\underline{z}}$ (3.14) at the level of $\overline{Q}$, we need an analogue of Van den Bergh’s outer double brackets. Taking an arbitrary $A$ as above, A right double bracket [20] is a bilinear map {{,}}$\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}:A\times A\to A\otimes A$ such that for $f,g,h\in A$,
@@
Let $Q^{\prime}$ be an arbitrary quiver with a total ordering $<^{\prime}$ on its arrows. For any $a,b\in Q^{\prime}$ with $a<^{\prime}b$, fix $y_{ab}\in\mathbb{C}$ and extend these constants to a skewsymmetric matrix $(y_{ab})_{a,b\in Q^{\prime}}$ by setting $y_{aa}=0$ and $y_{ab}=-y_{ba}$ if $b<^{\prime}a$.
@@
    Given a right double bracket {{,}}$\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}$ on an algebra $A$, introduce the multilinear map
    {{,,}}$\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}$ $:A^{\times 3}\to A^{\otimes 3}\,,$ (4.6)
    {{,,}}$\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}$ $:=\sum_{k=0,1,2}\tau_{(123)}^{k}\circ(\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}\otimes\operatorname{id}_{A})\circ(\operatorname{id}_{A}\otimes\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r})\circ\tau_{(123)}^{-k}\,,$
@@
    {{,,}}$\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}-\tau_{(12)}^{-1}\circ\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}\circ\tau_{(12)}=0\,.$ (4.7)
@@
For (2), we use that the map {{,,}}$\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}-\tau_{(12)}^{-1}\circ\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}\circ\tau_{(12)}$ is a derivation in each argument (for specific $A$-bimodule structures on $A^{\otimes 3}$) by [20, Lem. 3.12]. Hence we only have to verify (4.7) on generators $a,b,c\in Q^{\prime}$, cf. [20, Lem. 2.11]. We first compute
({{,}}$(\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}\otimes\operatorname{id}_{A})\circ(\operatorname{id}_{A}\otimes\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r})(a,b,c)=y_{bc}\left\{\!\!\left\{a,b\right\}\!\!\right\}_{r}\otimes c=y_{ab}y_{bc}\,a\otimes b\otimes c\,.$ (4.8)
@@
Indeed, this corresponds to taking the constants $y_{ab}=\epsilon(a)\epsilon(b)z_{ab}$. Thanks to [20, Prop. 3.15], the $(12)$-weak Poisson property guarantees that $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with a Poisson bracket, denoted {,}$\{-,-\}_{\underline{z}}$, uniquely determined by
@@
The morphism of algebras $\mathtt{X}_{\operatorname{tr}}:\operatorname{Sym}(H_{0}(A_{Q}))\to\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ (4.4) is a morphism of Poisson algebras if $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with {,}$\{-,-\}_{\underline{z}}$ (4.13) and $\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ is equipped with $\psi_{\underline{z}}$ (3.13).
@@
$\{(\mathtt{X}_{a})_{ij},(\mathtt{X}_{b})_{kl}\}=\left\{\begin{array}[]{ll}z_{ab}\,(\mathtt{X}_{a})_{ij}(\mathtt{X}_{b})_{kl}&\text{if }a,b\in Q\text{ with }a<b,\\ z_{a^{\ast}b^{\ast}}\,(\mathtt{X}_{a})_{ij}(\mathtt{X}_{b})_{kl}&\text{if }a^{\ast},b^{\ast}\in Q\text{ with }a^{\ast}<b^{\ast},\\ -z_{a^{\ast}b}\,(\mathtt{X}_{a})_{ij}(\mathtt{X}_{b})_{kl}&\text{if }a^{\ast},b\in Q\text{ with }a^{\ast}<b,\\ -z_{ab^{\ast}}\,(\mathtt{X}_{a})_{ij}(\mathtt{X}_{b})_{kl}&\text{if }a,b^{\ast}\in Q\text{ with }a<b^{\ast}.\end{array}\right.$ (4.15)
By definition of the skewsymmetric matrix $(z_{ab})_{a,b\in\overline{Q}}$ using (4.11), this is equivalent to (4.14). ∎
@@
If we combine Propositions 4.1 and 4.4, the only part of the statement which is unclear is that the linear combination $z_{0}\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}}$ is a Poisson bracket. Since we know that it is true if $z_{0}=0$ or $\underline{z}=0$, it suffices to show by rescaling that {,}$\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}}$ is a Poisson bracket on $\operatorname{Sym}(H_{0}(A_{Q}))$. Using again the known cases $z_{0}=0$ and $\underline{z}=0$, we are left to check that
[…]
@@
Let us define similarly $\mathtt{c}_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\{f,g,h\}$, noting that $\mathtt{c}_{f_{1},f_{2}}=-\mathtt{c}_{f_{2},f_{1}}$ by skewsymmetry of $(z_{ab})_{a,b\in\overline{Q}}$. We obtain by the derivation rule of the Poisson bracket {,}$\{-,-\}_{\mathtt{VdB}}$ that
@@
[…]
@@
$\mathcal{F}_{i,j}:=$ $\operatorname{tr}(\left\{\!\!\left\{a_{i},b_{j}\right\}\!\!\right\}_{o}^{\prime}a_{i+1}\cdots a_{n}a_{1}\cdots a_{i-1}\left\{\!\!\left\{a_{i},b_{j}\right\}\!\!\right\}_{o}^{\prime\prime}b_{j+1}\cdots b_{l}b_{1}\cdots b_{j-1})\,.$
@@
this expression vanishes since $(z_{ab})$ is symmetric in each index under the involution $a\mapsto a^{\ast}$ by (4.11). Therefore, $\hat{c}\,\mathcal{F}_{i,j}^{(1)}=(-\mathtt{c}_{g,h}+\mathtt{c}_{h,f})\mathcal{F}_{i,j}^{(1)}$, so that $\{h,\mathcal{F}_{i,j}^{(r)}\}_{\underline{z}}=(-\mathtt{c}_{g,h}+\mathtt{c}_{h,f})\,h\,\mathcal{F}_{i,j}^{(r)}$ for each $1\leq r\leq 7$. In particular, $\{h,\mathcal{F}_{i,j}\}_{\underline{z}}=(-\mathtt{c}_{g,h}+\mathtt{c}_{h,f})\,h\,\mathcal{F}_{i,j}$. Combining these observations with (4.20), we find

### v3

@@ 3.3.1. Deformation of the moment map
$P^{\gamma}$= $\,\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\gamma_{a}\operatorname{Id}+\mathtt{X}_{a^{\ast}}\mathtt{X}_{a})\,\partial_{a}\wedge\partial_{a^{\ast}}\right]$ (3.16)
@@
$\omega^{\gamma}$= $\,-\frac{1}{2}\sum_{a\in\overline{Q}}\epsilon(a)\,\operatorname{tr}\left[(\gamma_{a}\operatorname{Id}+\mathtt{X}_{a}\mathtt{X}_{a^{\ast}})^{-1}\,\mathrm{d}\mathtt{X}_{a}\wedge\mathrm{d}\mathtt{X}_{a^{\ast}}\right]$ (3.17)
@@ 3.3.2. Character varieties
The smooth complex variety $M_{g,r,n}$ is endowed with a Hamiltonian quasi-Poisson pencil of order g(g1)/2g(g-1)/2 centered at the quasi-Poisson bivector
$P_{\mathrm{char}}$= $\,\frac{1}{2}\sum_{1\leq i\leq g}\,\operatorname{tr}\left[A_{i}^{\ast}A_{i}\,\partial_{A_{i}}\wedge\partial_{A_{i}^{\ast}}-A_{i}A_{i}^{\ast}\partial_{A_{i}^{\ast}}\wedge\partial_{A_{i}}\right]+\frac{1}{2}\sum_{1\leq j\leq r}\operatorname{tr}\left[Z_{j}^{2}\,\partial_{Z_{j}}\wedge\partial_{Z_{j}}\right]$ (3.19)
@@ 3.3.3. Generalized MQV
$\Phi^{\mathfrak{c}}:M_{\overline{Q}}^{\mathfrak{c}}\to\operatorname{GL}_{\mathbf{n}},\quad\Phi^{\mathfrak{c}}(\mathtt{X})=(\Phi^{\mathfrak{c}}_{s}(\mathtt{X}))_{s\in I}\quad\text{ for}\quad\Phi^{\mathfrak{c}}_{s}(\mathtt{X}):=\prod_{\begin{subarray}{c}c\in C\\ \text{with}\,s\in I_{c}\end{subarray}}\Phi_{c,s}\in\operatorname{GL}_{n_{s}}.$ (3.24)
@@
$\mathrm{Inf}^{(c,r)}:=\sum_{\begin{subarray}{c}a\in\overline{Q}\,\text{with}\\ \mathfrak{c}(a)=c,\,t(a)=r\end{subarray}}\operatorname{tr}\left[\mathtt{X}_{a^{\ast}}\,\partial_{a^{\ast}}-\mathtt{X}_{a}\,\partial_{a}\right]\,.$ (3.27)
@@ 3.3.4. Quiver varieties
It was suggested by A. Alekseev that the Poisson structures defined in the previous two results can be obtained as special instances of a construction due to Lu and Mouquin [30]. Namely, for Proposition 3.19, see $G$ as a Poisson-Lie group for the zero Poisson structure, so that $\mathfrak{g}=\mathfrak{g}^{\prime}\oplus\bigoplus_{j=1}^{\ell}\mathfrak{g}_{j}$ is a Lie bialgebra with cobracket $\delta_{\mathfrak{g}}$ being trivially the sum of the zero cobrackets on $\mathfrak{g}^{\prime},\mathfrak{g}_{1},\ldots,\mathfrak{g}_{\ell}$. If $1^{(j)}\in\mathfrak{g}$ denotes the image of $1\in\mathbb{C}$ induced at the level of Lie algebras by the inclusions $\mathbb{C}^{\times}\hookrightarrow G_{j}\hookrightarrow G$, then $\mathtt{t}=\sum_{i<j}z_{ij}1^{(i)}\wedge 1^{(j)}\in\wedge^{2}\mathfrak{g}$ is a mixed twisting element [30, Def. 3.1]. It follows from [30, Lem. 2.6] that $(M,P+\mathtt{t}_{M})$ is a Poisson $(\mathfrak{g},\delta_{\mathfrak{g},\mathtt{t}})$-variety for the twisting cobracket $\delta_{\mathfrak{g},\mathtt{t}}=\delta_{\mathfrak{g}}+[\mathtt{t},-]$; but the latter vanishes while the action of $\mathfrak{g}$ on $M$ sends $\mathtt{t}$ to $\mathtt{t}_{M}=\psi_{\underline{z}}$ (3.4), hence $(M,P+\psi_{\underline{z}})$ is a Poisson $G$-variety.
As for Theorem 3.20, consider for each $a\in Q$ the subquiver $\Gamma_{a}=t(a)\underset{a^{\ast}}{\stackrel{{\scriptstyle a}}{{\rightleftarrows}}}h(a)$ of $\overline{Q}$. We have an action of $G_{a}:=\operatorname{GL}_{n_{t(a)}}\times\operatorname{GL}_{n_{h(a)}}$ on $Y_{a}:=M_{\Gamma_{a}}(n_{t(a)},n_{h(a)})$, cf. (2.10), thus $M_{\overline{Q}}=\times_{a\in Q}Y_{a}$ inherits an action of $\times_{a\in Q}G_{a}$ (it is “before fusion”, see Remark 2.11). Denote the Lie algebra of $G_{a}$ by $\mathfrak{g}_{a}$, and see $\mathfrak{g}=\oplus_{a\in Q}\mathfrak{g}_{a}$ as a Lie bialgebra for the zero cobracket as above. Let $1^{(a)}$ be the image of $(\operatorname{Id}_{n_{t(a)}},0_{n_{h(a)}})$ under the inclusion $\mathfrak{g}_{a}\hookrightarrow\mathfrak{g}$. Again, $\mathtt{t}=\sum_{a<b}z_{ab}1^{(a)}\wedge 1^{(b)}$ is mixed twisting and induces $\psi_{\underline{z}}$ (3.13). Hence $P_{\mathtt{qv}}+\psi_{\underline{z}}$ is a mixed product Poisson structure on $M_{\overline{Q}}$ for the action of $\times_{a\in Q}G_{a}$, cf. [30, Prop. 3.4]. Factoring the action of $\operatorname{GL}_{\mathbf{n}}$ through $\times_{a\in Q}G_{a}$ by taking diagonal embeddings yields the Poisson $\operatorname{GL}_{\mathbf{n}}$-variety structure considered in Theorem 3.20.
@@ 4. Noncommutative origin of the pencil
Fix an associative unital $\mathbb{C}$-algebra $A$, and consider $A\otimes A$ as a $\mathbb{C}$-algebra for the product satisfying $(f_{1}\otimes f_{2})(f_{3}\otimes f_{4})=f_{1}f_{3}\otimes f_{2}f_{4}$, $f_{1},\ldots,f_{4}\in A$. Denote by $\tau_{(12)}\in\operatorname{End}(A\otimes A)$ the operation that exchanges the two tensor factors. An outer double bracket [^9] This is simply called a double bracket by Van den Bergh [39]. We dub this operation as outer because the derivation rule of the second argument is defined thanks to the outer $A$-bimodule structure on $A\otimes A$ which is determined by $f_{1}\ast(g\otimes h)\ast f_{2}=f_{1}g\otimes hf_{2}$ for any $f_{1},f_{2},g,h\in A$. Below, we shall use right double brackets [20] based on the right $A$-bimodule structure $f_{1}\cdot_{r}(g\otimes h)\cdot_{r}f_{2}=g\otimes f_{1}hf_{2}$; hence qualifying the type of double bracket removes any possible confusion. is a bilinear map $\left\{\!\!\left\{-,-\right\}\!\!\right\}_{o}:A\times A\to A\otimes A$ such that for $f,g,h\in A$,
@@
An important result is that the pair $(A_{Q},\left\{\!\!\left\{-,-\right\}\!\!\right\}_{o})$ is a double quasi-Poisson algebra [15, Theorem 3.3]. Without unravelling the definition (cf. [39, §5.1]) let us only give consequences of this fact. Put $H_{0}(A_{Q})=A_{Q}/[A_{Q},A_{Q}]$, which is a vector space coming with a linear map $\operatorname{tr}:A_{Q}\to H_{0}(A_{Q})$ that sends commutators to zero. We get from [39, Lem. 5.1.3] that $H_{0}(A_{Q})$ is equipped with a Lie bracket, denoted $\{-,-\}_{\mathtt{VdB}}$, uniquely determined by
@@
where $\mathrm{m}:A_{Q}\otimes A_{Q}\to A_{Q}$ is the multiplication. In particular, (4.3) extends to a Poisson bracket on $\operatorname{Sym}(H_{0}(A_{Q}))$, which we also denote by $\{-,-\}_{\mathtt{VdB}}$. The following result follows from [39, §7.7,7.13] and the fact that the representation space $\operatorname{Rep}(A_{Q},\mathbf{n})$ coincides with $M_{\overline{Q}}^{\gamma}(\mathbf{n})$ for any $\mathbf{n}\in\mathbb{Z}^{I}_{\geq 0}$.
@@
is a morphism of Poisson algebras if $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with $\{-,-\}_{\mathtt{VdB}}$ (4.3) and $\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ is equipped with $P^{\gamma}$ (3.16).
@@
To understand the bivector $\psi_{\underline{z}}$ (3.14) at the level of $\overline{Q}$, we need an analogue of Van den Bergh’s outer double brackets. Taking an arbitrary $A$ as above, A right double bracket [20] is a bilinear map $\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}:A\times A\to A\otimes A$ such that for $f,g,h\in A$,
@@
Let $Q^{\prime}$ be an arbitrary quiver with a total ordering $^{\prime}$ on its arrows. For any $a,b\in Q^{\prime}$ with $a<^{\prime}b$, fix $y_{ab}\in\mathbb{C}$ and extend these constants to a skewsymmetric matrix $(y_{ab})_{a,b\in Q^{\prime}}$ by setting $y_{aa}=0$ and $y_{ab}=-y_{ba}$ if $b<^{\prime}a$.
@@
    Given a right double bracket $\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}$ on an algebra $A$, introduce the multilinear map
    $\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}$ $:A^{\times 3}\to A^{\otimes 3}\,$, (4.6)
    $\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}$ $=\sum_{k=0,1,2}\tau_{(123)}^{k}\circ(\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}\otimes\operatorname{id}_{A})\circ(\operatorname{id}_{A}\otimes\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r})\circ\tau_{(123)}^{-k}\,$,
@@
    $\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}-\tau_{(12)}^{-1}\circ\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}\circ\tau_{(12)}=0\,.$ (4.7)
@@
For (2), we use that the map $\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}-\tau_{(12)}^{-1}\circ\left\{\!\!\left\{-,-,-\right\}\!\!\right\}_{r}\circ\tau_{(12)}$ is a derivation in each argument (for specific $A$-bimodule structures on $A^{\otimes 3}$) by [20, Lem. 3.12]. Hence we only have to verify (4.7) on generators $a,b,c\in Q^{\prime}$, cf. [20, Lem. 2.11]. We first compute
$(\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r}\otimes\operatorname{id}_{A})\circ(\operatorname{id}_{A}\otimes\left\{\!\!\left\{-,-\right\}\!\!\right\}_{r})(a,b,c)=y_{bc}\left\{\!\!\left\{a,b\right\}\!\!\right\}_{r}\otimes c=y_{ab}y_{bc}\,a\otimes b\otimes c\,.$ (4.8)
@@
Indeed, this corresponds to taking the constants $y_{ab}=\epsilon(a)\epsilon(b)z_{ab}$. Thanks to [20, Prop. 3.15], the $(12)$-weak Poisson property guarantees that $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with a Poisson bracket, denoted $\{-,-\}_{\underline{z}}$, uniquely determined by
@@
The morphism of algebras $\mathtt{X}_{\operatorname{tr}}:\operatorname{Sym}(H_{0}(A_{Q}))\to\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ (4.4) is a morphism of Poisson algebras if $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with $\{-,-\}_{\underline{z}}$ (4.13) and $\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ is equipped with $\psi_{\underline{z}}$ (3.13).
@@
By [20, Theorem 4.11] and (4.12), the operation $\{-,-\}$ is a Poisson bracket, and it descends to $\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$. We remark in view of [20, (4.7)] that $\mathtt{X}_{\operatorname{tr}}$ is a morphism of Poisson algebras if $\mathbb{C}[M_{\overline{Q}}^{\gamma}/\!/\operatorname{GL}_{\mathbf{n}}]$ is equipped with that Poisson bracket and $\operatorname{Sym}(H_{0}(A_{Q}))$ is equipped with $\{-,-\}_{\underline{z}}$ (4.13). Thus, we need to verify that (4.14) coincides with the expression for $\psi_{\underline{z}}$ (3.13) evaluated on generators $(\mathtt{X}_{a})_{ij}$ ($a\in\overline{Q}$) of $\mathbb{C}[M_{\overline{Q}}^{\gamma}]$. Using (3.13), we compute
@@
By definition of the skewsymmetric matrix $(z_{ab})_{a,b\in\overline{Q}}$ using (LABEL:Eq:zab-Ext), this is equivalent to (4.14). ∎
@@
If we combine Propositions 4.1 and 4.4, the only part of the statement which is unclear is that the linear combination $z_{0}\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}}$ is a Poisson bracket. Since we know that it is true if $z_{0}=0$ or $\underline{z}=0$, it suffices to show by rescaling that $\{-,-\}_{\mathtt{VdB}}+\{-,-\}_{\underline{z}}$ is a Poisson bracket on $\operatorname{Sym}(H_{0}(A_{Q}))$. Using again the known cases $z_{0}=0$ and $\underline{z}=0$, we are left to check that
$\{f,\{g,h\}_{\mathtt{VdB}}\}_{\underline{z}}+\{f,\{g,h\}_{\underline{z}}\}_{\mathtt{VdB}}+\text{cyclic perm. }=0\,$, (4.16)
@@
Let us define similarly $\mathtt{c}_{f_{1},f_{2}}$ for any $f_{1},f_{2}\in\{f,g,h\}$, noting that $\mathtt{c}_{f_{1},f_{2}}=-\mathtt{c}_{f_{2},f_{1}}$ by skewsymmetry of $(z_{ab})_{a,b\in\overline{Q}}$. We obtain by the derivation rule of the Poisson bracket $\{-,-\}_{\mathtt{VdB}}$ that
@@
$\{h,\{f,g\}_{\mathtt{VdB}}\}_{\underline{z}}+\mathtt{c}_{g,h}\{f,g\}_{\mathtt{VdB}}\,h+\mathtt{c}_{h,f}\{g,f\}_{\mathtt{VdB}}\,h+\text{cyclic perm. }=0\,.$ (4.18)
@@
$\mathcal{F}_{i,j}$= $\operatorname{tr}(\left\{\!\!\left\{a_{i},b_{j}\right\}\!\!\right\}_{o}^{\prime}a_{i+1}\cdots a_{n}a_{1}\cdots a_{i-1}\left\{\!\!\left\{a_{i},b_{j}\right\}\!\!\right\}_{o}^{\prime\prime}b_{j+1}\cdots b_{l}b_{1}\cdots b_{j-1})\,.$
@@
this expression vanishes since $(z_{ab})$ is symmetric in each index under the involution $a\mapsto a^{\ast}$ by (LABEL:Eq:zab-Ext). Therefore, $\hat{c}\,\mathcal{F}_{i,j}^{(1)}=(-\mathtt{c}_{g,h}+\mathtt{c}_{h,f})\mathcal{F}_{i,j}^{(1)}$, so that $\{h,\mathcal{F}_{i,j}^{(r)}\}_{\underline{z}}=(-\mathtt{c}_{g,h}+\mathtt{c}_{h,f})\,h\,\mathcal{F}_{i,j}^{(r)}$ for each $1\leq r\leq 7$. In particular, $\{h,\mathcal{F}_{i,j}\}_{\underline{z}}=(-\mathtt{c}_{g,h}+\mathtt{c}_{h,f})\,h\,\mathcal{F}_{i,j}$. Combining these observations with (4.20), we find